# Statistical Guarantees of Correctness Coverage for Medical Multiple-Choice Question Answering

### v1

@@ 4.1 Experimental Settings
In this study, we utilized three widely recognized benchmark datasets—MedMCQA, MedQA, and the multi-task MMLU dataset—to rigorously assess the performance of our proposed approach across diverse medical question-answering (QA) scenarios. To ensure a robust and unbiased evaluation, we uniformly sampled 2,000 representative instances from each dataset, maintaining consistency and comparability throughout our experimental design.
@@
The selected datasets provide a broad and diverse benchmark for evaluating medical and general knowledge QA systems. The MedMCQA dataset is specifically curated for real-world medical entrance examinations, encompassing over 194,000 high-quality multiple-choice questions. It spans 2,400 healthcare topics across 21 medical subjects, requiring advanced reasoning and analytical skills. The MedQA dataset offers a multilingual, free-form multiple-choice question set. It includes 12,723 questions in English, 34,251 in simplified Chinese, and 14,123 in traditional Chinese, along with a large-scale corpus of medical textbook content to support reading comprehension models.
@@
The MMLU dataset evaluates multi-task language understanding across 57 diverse tasks, including humanities, social sciences, hard sciences, and law. It challenges models with both general knowledge queries and advanced problem-solving scenarios. Together, these datasets serve as robust and diverse evaluation tools, enabling a thorough assessment of model performance in both specialized medical and broad general knowledge domains, and facilitating the advancement of more reliable and adaptable QA systems.
Base LLMs
@@
In this study, we assess the performance of our proposed framework using four distinct large language models (LLMs) with varying architectures and parameter scales. Specifically, the models employed include Llama-3.2-1B-Instruct, Llama-3.2-3B-Instruct, Qwen2.5-1.5B-Instruct, and Qwen2.5-3B-Instruct. These models span a parameter range from 1.5 billion to 3 billion, enabling a comprehensive analysis of how model size influences performance across diverse tasks.
@@
The Llama-3.2-1B-Instruct and Llama-3.2-3B-Instruct models are built upon the Llama architecture, which is explicitly designed to excel in instruction-following scenarios. These models are particularly effective in tasks requiring a nuanced understanding of natural language instructions and the generation of coherent and contextually appropriate responses. Conversely, the Qwen2.5-1.5B-Instruct and Qwen2.5-3B-Instruct models are optimized for multi-task learning, demonstrating versatility across a wide range of natural language processing (NLP) tasks, including question answering, classification, and reasoning. By incorporating LLMs with diverse architectures and parameter scales, our experimental design enhances the robustness and generalizability of the evaluation. This approach ensures a thorough and nuanced assessment of our framework’s effectiveness across models of varying sizes and functional capabilities, contributing to a more holistic understanding of its performance dynamics.
Hyperparameters
@@
In our experiments, the split ratio between the calibration and test sets is fixed at 0.5, ensuring an equal distribution of data across both sets. This strategy provides a balanced and reliable evaluation of model performance. Consistent with prior studies, we utilize multinomial sampling to generate $M$ candidate responses for each data point. For the MMLU and MedMCQA datasets, which comprise four answer options per question, the number of candidate responses $M$ is set to 20, aligning with established practices [21]. For the MMLU-Pro dataset, where each sample includes ten multiple-choice options, $M$ is increased to 50 to more accurately approximate the model output distribution.
@@
To optimize computational efficiency in multiple-choice question-answering (MCQA) tasks, we set the maximum generation length to 1 and the generation temperature to 1.0 for sampling. For open-domain QA tasks, a broader set of hyperparameters is explored, with further performance refinement achieved through risk calibration techniques as detailed in prior research [2, 18, 15]. The parameter input_length refers to the embedding length of the input prompt after being encoded by the tokenizer of the current language model. Additionally, for both tasks, the maximum generation length is set to 36 tokens, ensuring that generated responses are sufficiently detailed while avoiding excessive constraints imposed by model limitations.
Evaluation Metrics
@@
The performance of our framework is rigorously assessed using two principal metrics: the Empirical Miscoverage Rate (EMR) and the Average Prediction Set Size (APSS). The EMR metric evaluates the degree to which our framework maintains the intended marginal coverage within the test set. It measures the frequency at which the true label of a test sample falls outside the predicted coverage set, offering a quantitative assessment of prediction reliability. By thoroughly analyzing the EMR, we verify that the model consistently achieves the desired coverage level under various test conditions, thereby demonstrating its robustness and generalization capabilities.
@@
In addition to EMR, the APSS metric is utilized to assess both prediction efficiency and uncertainty for each LLM. The APSS represents the average number of candidate responses generated per test sample, serving as an indicator of the model’s handling of uncertainty in the prediction process. A lower APSS reflects a more efficient model that produces fewer, yet accurate, responses, while a higher APSS may indicate increased uncertainty or a more conservative predictive approach. By combining EMR and APSS metrics, our evaluation framework provides a comprehensive analysis of model performance, balancing predictive reliability with efficiency and delivering valuable insights into the model’s behavior in real-world applications.
@@
Figure 3: EMR performance on the MedQA dataset, demonstrating the consistent reliability of Llama models over a range of error rates, with the Qwen2.5 models showing more variability.
4.2 Empirical Evaluations
Statistical Guarantees of the EMR Metric
This section presents a rigorous validation of the calibrated prediction sets constructed according to Eq. 2, demonstrating that they provide robust correctness coverage guarantees under various user-specified error rates. Additionally, we examine the practical utility of these prediction sets through selective prediction based on our proposed uncertainty measure.
Empirical Coverage Guarantees
To empirically validate the lower bound of the correctness coverage rate, we partition each of the four datasets using a 1:10 split, allocating the subsets as calibration and test sets, respectively. The calibration set is employed to establish the conformal uncertainty criterion, guided by the upper bound of the permissible error rate. The correctness coverage rate is subsequently evaluated on the test set, with the results for the MedMCQA, MedQA, and MMLU datasets illustrated in Figures 2, 3, and 4, respectively.
In Figure 2, which displays the MedMCQA dataset results, a consistent trend of increasing error rates is observed as the confidence level $\alpha$ rises. Notably, the Llama-3.2-1B-Instruct and Llama-3.2-3B-Instruct models consistently maintain lower and more stable error rates compared to the Qwen2.5-1.5B-Instruct and Qwen2.5-3B-Instruct models, which exhibit greater variability.
A similar pattern is evident in Figure 3, representing the MedQA dataset. The Llama models again demonstrate superior stability and lower error rates, whereas the Qwen2.5 models show an increase in error rates at higher $\alpha$ values. This finding indicates that the Llama models offer more reliable predictive performance across a broad range of confidence levels.
@@ 4.2 Empirical Evaluations
For the MMLU dataset, as shown in Figure 4, we analyze four distinct categories: high school biology, anatomy, clinical knowledge, and college medicine. Across all categories, the Llama models consistently exhibit enhanced stability and lower error rates relative to the Qwen2.5 models. The performance discrepancy is most pronounced in the clinical knowledge category, where the Qwen2.5 models show significant fluctuations in error rates.
Overall, these empirical results validate the stringent control over the correctness coverage rate under varying error rate conditions across all datasets. The statistical guarantees provided by these findings align closely with the theoretical framework established in Eqs. 2 and 7 demonstrating that the calibrated prediction sets constructed by our method deliver robust and statistically rigorous correctness coverage guarantees under diverse calibration scenarios.
Uncertainty Estimation of LLMs
The Average Prediction Set Size (APSS) metric serves as a critical indicator of model uncertainty, demonstrating a clear inverse relationship with the risk level $\alpha$. As illustrated in Table 1, the APSS consistently decreases as the risk level increases, reflecting a more stringent criterion for constructing prediction sets. This trend is evident across all evaluated models, underscoring the robustness of the APSS metric as a measure of predictive uncertainty.
For each dataset, an increase in the risk level $\alpha$ generally results in a reduction of the APSS metric. Larger models, such as Llama-3.2-1B-Instruct and Llama-3.2-3B-Instruct, typically exhibit smaller average prediction set sizes compared to the Qwen2.5 series models. For instance, in the MedMCQA dataset, the Llama-3.2-1B-Instruct model reduces its APSS from 3.569 at $\alpha=0.1$ to 0.214 at $\alpha=0.9$. Similarly, in the MedQA dataset, the Llama-3.2-3B-Instruct model demonstrates a pronounced decrease in APSS from 4.000 at $\alpha=0.1$ to 0.137 at $\alpha=0.9$, highlighting a consistent narrowing of the prediction sets as the uncertainty threshold tightens.
In contrast, the Qwen2.5 models exhibit a more gradual reduction in APSS with increasing $\alpha$ values. Specifically, the Qwen2.5-1.5B-Instruct model on the MMLU dataset maintains an APSS of 4.000 up to $\alpha=0.5$, after which a gradual decrease is observed. This behavior suggests that the Qwen2.5 models tend to generate larger and more stable prediction sets across varying confidence levels, potentially indicating higher uncertainty in their predictions.
@@
Overall, the APSS metric not only serves as a reliable measure of model uncertainty but also offers critical insights into model optimization strategies for achieving more precise and selective predictions under varying confidence thresholds. This reinforces the utility of APSS as a tool for evaluating model performance in high-stakes applications where prediction certainty is paramount.
Table 1: Results of the APSS metric at various risk levels.
Datasets LLMs / $\alpha$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 MedMCQA Llama-3.2-1B-Instruct 3.57 3.15 2.58 2.03 1.51 1.08 1.08 0.50 0.22 Llama-3.2-3B-Instruct 4.00 2.60 1.91 1.37 1.01 0.75 0.48 0.28 0.12 Qwen2.5-1.5B-Instruct 4.00 4.00 4.00 1.98 1.54 1.11 0.74 0.44 0.33 Qwen2.5-3B-Instruct 4.00 4.00 4.00 1.45 1.17 0.75 0.45 0.45 0.45 MedQA Llama-3.2-1B-Instruct 4.00 4.00 3.35 2.74 2.12 1.57 1.09 0.77 0.24 Llama-3.2-3B-Instruct 4.00 4.00 1.94 1.42 1.06 0.67 0.44 0.29 0.21 Qwen2.5-1.5B-Instruct 4.00 4.00 4.00 4.00 4.00 1.30 0.83 0.47 0.38 Qwen2.5-3B-Instruct 4.00 4.00 4.00 4.00 4.00 0.91 0.64 0.64 0.64 MMLU Llama-3.2-1B-Instruct 3.89 3.32 2.83 2.14 1.53 1.04 0.68 0.21 0.03 Llama-3.2-3B-Instruct 2.43 1.66 1.24 0.92 0.65 0.49 0.25 0.25 0.25 Qwen2.5-1.5B-Instruct 4.00 4.00 1.41 1.06 0.79 0.54 0.43 0.43 0.43 Qwen2.5-3B-Instruct 4.00 4.00 1.06 0.83 0.83 0.83 0.83 0.83 0.83
@@ 4.3 Sensitivity Analysis
EMR at Various Split Ratios
To assess the sensitivity of our method to different data partitioning strategies, we conducted a comprehensive analysis of the Exact Match Ratio (EMR) on the MedMCQA dataset, with the parameter $\alpha=0.1$. The evaluation was performed across a range of split ratios (0.1, 0.3, 0.5, 0.7), as detailed in Table 2. The results demonstrate that the Llama-3.2-1B-Instruct and Llama-3.2-3B-Instruct models maintained consistently low and stable EMR values, ranging from 0.13 to 0.20, irrespective of the split ratio. In contrast, the Qwen2.5-1.5B-Instruct and Qwen2.5-3B-Instruct models exhibited an EMR consistently at 0.00 across all tested scenarios, indicating robust performance independent of data partitioning.
These findings underscore the robustness of our proposed method, demonstrating its capacity to sustain stable performance across diverse data distributions and varying sample sizes. This stability is particularly valuable in mitigating the risks associated with model overfitting or underfitting, thereby enhancing the method’s generalization capabilities across different experimental conditions.
Table 2: EMR at various split ratios.
Datasets LLMs / $\alpha=0.1$ split ratios=0.1 0.3 0.5 0.7 MedMCQA Llama-3.2-1B-Instruct 0.14 0.14 0.14 0.14 Llama-3.2-3B-Instruct 0.13 0.13 0.20 0.20 Qwen2.5-1.5B-Instruct 0.00 0.00 0.00 0.00 Qwen2.5-3B-Instruct 0.00 0.00 0.00 0.00
AUROC Analysis

### v2

@@ 4 Experiments
The experiments presented in this section aim to evaluate the performance of the proposed framework, utilizing a set of widely recognized medical datasets and state-of-the-art large language models (LLMs). These experiments focus on assessing the framework’s ability to manage uncertainty quantification (UQ) and task-specific risks across various medical question-answering (QA) tasks. Additionally, the performance of the framework is measured using several standard metrics, providing a comprehensive analysis of its effectiveness in real-world applications.
@@ 4.1.1 Datasets
The study utilizes three widely recognized benchmark datasets—MedMCQA, MedQA, and the multi-task MMLU dataset—to rigorously assess the performance of the proposed approach across diverse medical question-answering (QA) scenarios. To ensure a robust and unbiased evaluation, we uniformly sampled 2,000 representative instances from each dataset to maintain consistency and comparability throughout the experimental design.
@@
The selected datasets provide a broad and diverse benchmark for evaluating medical and general knowledge QA systems. The MedMCQA dataset is curated explicitly for real-world medical entrance examinations, encompassing over 194,000 high-quality multiple-choice questions [35]. It spans 2,400 healthcare topics across 21 medical subjects, requiring advanced reasoning and analytical skills. The MedQA dataset offers a multilingual multiple-choice question dataset. It includes 12,723 questions in English, 34,251 in simplified Chinese, and 14,123 in traditional Chinese, along with a large-scale corpus of medical textbook content to support reading comprehension models.
@@
The MMLU dataset evaluates multi-task language understanding across 57 diverse tasks, including humanities, social sciences, hard sciences, and law. It challenges models with both general knowledge queries and advanced problem-solving scenarios. Together, these datasets serve as robust and diverse evaluation tools, enabling a thorough assessment of model performance in specialized medical and broad general knowledge domains, and facilitating the advancement of more reliable and adaptable QA systems.
4.1.2 Based LLMs
@@
This work evaluates the proposed framework using four large language models (LLMs), each with distinct architectures and model sizes. Specifically, the models employed include Llama-3.2-1B-Instruct [43], Llama-3.2-3B-Instruct [43], Qwen2.5-1.5B-Instruct [44], and Qwen2.5-3B-Instruct [44]. These models span a parameter range from 1.5 billion to 3 billion, enabling a comprehensive analysis of how model size influences performance across diverse tasks.
@@
The Llama-3.2-1B-Instruct [43] and Llama-3.2-3B-Instruct [43] models are built upon the Llama architecture, which is explicitly designed to excel in instruction-following scenarios. These models are particularly effective in tasks requiring a nuanced understanding of natural language instructions and generating coherent and contextually appropriate responses. Conversely, the Qwen2.5-1.5B-Instruct [44] and Qwen2.5-3B-Instruct [44] models are optimized for multi-task learning, performing effectively in various NLP applications, including classification, reasoning, and responding to questions. By incorporating LLMs with diverse architectures and parameter scales, the proposed experimental design enhances the robustness and generalizability of the evaluation. This approach ensures a thorough and nuanced assessment of the proposed framework’s effectiveness across models of varying sizes and functional capabilities, contributing to a more holistic understanding of its performance dynamics.
4.1.3 Hyperparameters
@@
In the experiments, the split ratio between the calibration and test sets is fixed at 0.5, ensuring an equal data distribution across both sets. This strategy provides a balanced and reliable evaluation of model performance. Consistent with prior studies, multinomial sampling is utilized to generate $M$ candidate responses for each data point. For the MMLU and MedMCQA datasets, each comprising four answer options per question, the number of candidate responses $M$ is set to 20, aligning with established practices [21]. For the MMLU-Pro dataset, where each sample includes ten multiple-choice options, $M$ is increased to 50 to approximate the model output distribution more accurately.
@@
To optimize computational efficiency in multiple-choice question-answering (MCQA) tasks, the maximum generation length is set to 1, and the generation temperature is fixed at 1.0 for sampling. A broader set of hyperparameters is explored for open-domain QA tasks, with further performance refinement achieved through risk calibration techniques as detailed in prior research [29, 28, 33]. The parameter input_length refers to the embedding length of the input prompt after being encoded by the tokenizer of the current language model. Additionally, the maximum generation length for both tasks is set to 36 tokens, ensuring that generated responses are sufficiently detailed while avoiding excessive constraints imposed by model limitations.
4.1.4 Evaluation Metrics
@@
The performance is rigorously assessed using two principal metrics: the Empirical Miscoverage Rate (EMR) and the Average Prediction Set Size (APSS). The EMR assesses how well the proposed framework maintains the intended marginal coverage within the test set. It measures the frequency at which the true label of a test sample falls outside the predicted coverage set, offering a quantitative assessment of prediction reliability. A thorough analysis of the EMR verifies that the model consistently achieves the desired coverage level under various test conditions, demonstrating its robustness and generalization capabilities.
@@
In addition to EMR, the APSS metric is utilized to assess prediction efficiency and uncertainty for each LLM. The APSS represents the average number of candidate responses generated per test sample, indicating the model’s handling of uncertainty in the prediction process. A lower APSS reflects a more efficient model that produces fewer yet accurate responses, while a higher APSS may indicate increased uncertainty or a more conservative predictive approach. By combining EMR and APSS metrics, the proposed framework offers a comprehensive analysis of model performance, balancing predictive reliability with efficiency and providing valuable insights into its behavior in real-world applications.
4.2 Empirical Evaluations
4.2.1 Statistical Guarantees of the EMR Metric
This section rigorously validates the calibrated prediction sets constructed according to Eq. 2, showing that they reliably achieve the desired coverage levels across different user-defined error rates. Furthermore, the practical effectiveness is investigated by applying selective prediction guided by the proposed uncertainty metric.
4.2.2 Empirical Coverage Guarantees
To verify the minimum correctness coverage level empirically, each of the four datasets is partitioned with a 1:10 ratio into calibration and testing subsets. The calibration subset determines the conformal uncertainty threshold based on the maximum allowable error rate. Subsequently, coverage performance is assessed using the testing subset, and corresponding results for MedMCQA, MedQA, and MMLU are shown in Figures 2, 3, and 4.
@@ 4.2.2 Empirical Coverage Guarantees
Figure 3: EMR performance on the MedQA dataset, demonstrating the consistent reliability of Llama models over a range of error rates, with the Qwen2.5 models showing various variability.
In Figure 2, which displays the MedMCQA dataset results, a consistent trend of increasing error rates is observed as the confidence level $\alpha$ rises. Notably, the Llama-3.2-1B-Instruct [43] and Llama-3.2-3B-Instruct [43] models consistently maintain lower and more stable error rates compared to the Qwen2.5-1.5B-Instruct [44] and Qwen2.5-3B-Instruct [44] models, which exhibit more significant variability.
A similar pattern is evident in Figure 3, representing the MedQA dataset. The Llama models again demonstrate superior stability and lower error rates, whereas the Qwen2.5 models show an increase in error rates at higher $\alpha$ values. The findings indicates that the Llama models offer more reliable predictive performance across a broad range of confidence levels.
For the MMLU dataset, as shown in Figure 4, four distinct categories are analyzed: high school biology, anatomy, clinical knowledge, and college medicine. Across all categories, the Llama models consistently exhibit enhanced stability and lower error rates relative to the Qwen2.5 models. The performance discrepancy is mainly pronounced in the clinical knowledge category, where the Qwen2.5 models show significant fluctuations in error rates.
@@
These empirical results validate the stringent control over the correctness coverage rate under varying error rate conditions across all datasets. The statistical guarantees provided by these findings align closely with the theoretical framework established in Eqs. 2 and 7, demonstrating that the calibrated prediction sets constructed by the proposed method ensure robust and statistically rigorous correctness coverage under diverse calibration scenarios.
4.2.3 Uncertainty Estimation of LLMs
The Average Prediction Set Size (APSS) metric is a critical indicator of model uncertainty, demonstrating a clear inverse relationship with the risk level $\alpha$. As illustrated in Table 1, the APSS consistently decreases as the risk level increases, reflecting a more stringent criterion for constructing prediction sets. This trend is evident across all evaluated models, underscoring the robustness of the APSS metric as a measure of predictive uncertainty.
For each dataset, an increase in the risk level $\alpha$ generally reduces the APSS metric. The more larger models typically exhibit smaller average prediction set sizes than the Qwen2.5 series models. For instance, in the MedMCQA dataset, the Llama-3.2-1B-Instruct [43] model reduces its APSS from 3.569 at $\alpha=0.1$ to 0.214 at $\alpha=0.9$. Similarly, in the MedQA dataset, the Llama-3.2-3B-Instruct model demonstrates a pronounced decrease in APSS from 4.000 at $\alpha=0.1$ to 0.137 at $\alpha=0.9$, highlighting a consistent narrowing of the prediction sets as the uncertainty threshold tightens.
Table 1: Results of the APSS metric at various risk levels.
Datasets LLMs / $\alpha$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
MedMCQA Llama-3.2-1B-Instruct 3.57 3.15 2.58 2.03 1.51 1.08 1.08 0.50 0.22
Llama-3.2-3B-Instruct 4.00 2.60 1.91 1.37 1.01 0.75 0.48 0.28 0.12
Qwen2.5-1.5B-Instruct 4.00 4.00 4.00 1.98 1.54 1.11 0.74 0.44 0.33
Qwen2.5-3B-Instruct 4.00 4.00 4.00 1.45 1.17 0.75 0.45 0.45 0.45
MedQA Llama-3.2-1B-Instruct 4.00 4.00 3.35 2.74 2.12 1.57 1.09 0.77 0.24
Llama-3.2-3B-Instruct 4.00 4.00 1.94 1.42 1.06 0.67 0.44 0.29 0.21
Qwen2.5-1.5B-Instruct 4.00 4.00 4.00 4.00 4.00 1.30 0.83 0.47 0.38
Qwen2.5-3B-Instruct 4.00 4.00 4.00 4.00 4.00 0.91 0.64 0.64 0.64
MMLU Llama-3.2-1B-Instruct 3.89 3.32 2.83 2.14 1.53 1.04 0.68 0.21 0.03
Llama-3.2-3B-Instruct 2.43 1.66 1.24 0.92 0.65 0.49 0.25 0.25 0.25
Qwen2.5-1.5B-Instruct 4.00 4.00 1.41 1.06 0.79 0.54 0.43 0.43 0.43
Qwen2.5-3B-Instruct 4.00 4.00 1.06 0.83 0.83 0.83 0.83 0.83 0.83
In contrast, the Qwen2.5 models exhibit a more gradual reduction in APSS with increasing $\alpha$ values. Specifically, the Qwen2.5-1.5B-Instruct model [44] on the MMLU dataset maintains an APSS of 4.000 up to $\alpha=0.5$, after which a gradual decrease is observed. This behavior suggests that the Qwen2.5 models tend to generate larger and more stable prediction sets across varying confidence levels, potentially indicating higher uncertainty in their predictions.
@@ 4.2.3 Uncertainty Estimation of LLMs
Overall, the APSS metric serves as a reliable measure of model uncertainty and offers critical insights into model optimization strategies for achieving more precise and selective predictions under varying confidence thresholds. It reinforces the utility of APSS as a tool for evaluating model performance in high-stakes applications where prediction certainty is paramount.
@@ 4.3 Sensitivity Analysis
4.3.1 EMR at Various Split Ratios
To assess the sensitivity of the proposed method to different data partitioning strategies, a comprehensive analysis of the Empirical Miscoverage Rate (EMR) was conducted on the MedMCQA dataset, with the parameter $\alpha=0.1$. The evaluation was performed across a range of split ratios (0.1, 0.3, 0.5, 0.7), as detailed in Table 2. The results demonstrate that the Llama-3.2-1B-Instruct [43] and Llama-3.2-3B-Instruct [43] models maintained consistently low and stable EMR values, ranging from 0.13 to 0.20, irrespective of the split ratio. In contrast, the Qwen2.5-1.5B-Instruct [44] and Qwen2.5-3B-Instruct [44] models exhibited an EMR consistently at 0.00 across all tested scenarios, indicating robust performance independent of data partitioning.
These findings underscore the robustness of the proposed method, demonstrating its capacity to maintain stable performance across diverse data distributions and varying sample sizes. This stability is particularly valuable in mitigating the risks associated with model overfitting or underfitting, enhancing the method’s generalization capabilities across different experimental conditions.
Table 2: Exact Match Ratio (EMR) at various split ratios.
Datasets LLMs / $\alpha=0.1$ Split Ratio = 0.1 0.3 0.5 0.7
MedMCQA Llama-3.2-1B-Instruct 0.14 0.14 0.14 0.14
Llama-3.2-3B-Instruct 0.13 0.13 0.20 0.20
Qwen2.5-1.5B-Instruct 0.00 0.00 0.00 0.00
Qwen2.5-3B-Instruct 0.00 0.00 0.00 0.00
4.3.2 AUROC Analysis